# Foundation Models for Causal Inference via Prior-Data Fitted Networks

### v2

@@ Appendix A Extended related work
Foundation models are an emerging paradigm that has revolutionized machine learning for various data modalities, particularly language and vision tasks (Devlin, 2018; Lahat et al., 2024; Touvron et al., 2023b; a). The same paradigm is now being explored for tabular data – a modality that underpins the large majority of analyses in science and business (van Breugel & van der Schaar, 2024). Prior-data-fitted networks (PFNs) (Müller et al., 2022) constitute a powerful approach to training tabular foundation models. PFNs are large transformers trained on synthetic data to perform Bayesian inference through in-context learning. TabPFN (Hollmann et al., 2023; 2025) scaled this idea by pairing the transformer with a Bayesian neural network prior over structural causal models (SCMs) and demonstrating state-of-the-art performance on various tabular benchmarks. Subsequent work extended PFNs to time–series forecasting (Hoo et al., 2025) and analyzed their in-context learning abilities theoretically (Nagler, 2023). Critically, all existing PFNs are trained only for predictive tasks and do not target causal estimands; they therefore are not designed for causal inference of treatment effects, which is the goal of our paper.
@@
Causal inference, such as the estimation of average treatment effects, originates from fields like econometrics (Imbens & Angrist, 1994; Angrist, 1990), statistics (van der Laan & Rubin, 2006), and epidemiology (Robins, 1986; 1994). Machine learning methods have been proposed to estimate heterogeneous effects to support personalized decision-making. One line of work are frequentist methods, which often build on semiparametric theory (Robins et al., 1994; Robins, 1999), yielding model-agnostic estimators that are doubly robust and Neyman-orthogonal (van der Laan, 2006; Chernozhukov et al., 2018; Nie & Wager, 2021; Foster & Syrgkanis, 2023; Kennedy, 2023a). Another line of work builds upon specific machine learning methods/ architectures such as regression trees (Wager & Athey, 2018) or neural networks (Johansson et al., 2016; Shalit et al., 2017a; Shi et al., 2019) and adopts them to causal inference. Bayesian alternatives include Bayesian additive regression trees (Hahn et al., 2020) or Gaussian-process counterfactual regression (Alaa & van der Schaar, 2017). However, all of the existing estimators above must be retrained for every new dataset. In contrast, our CausalFM allows for pre-trained models to approximate Bayesian causal inference.
@@ Appendix C Proofs
[…]
$Q(\mathbb{P}_{\mathrm{int}}^{\mathcal{W}})\;=\;\bar{Q}(\mathbb{P}_{\mathrm{obs}}^{\mathcal{W}})\;=\;\bar{Q}(\mathbb{P}_{\mathrm{obs}}^{\mathcal{S}_{0}})\neq Q(\mathbb{P}_{\mathrm{int}}^{\mathcal{S}_{0}}).$ (12)
Draw data $\mathcal{D}_{n}\sim\mathbb{P}_{\mathrm{obs}}^{\mathcal{W}}$. For every $\mathcal{S}\in\mathcal{E}(\mathbb{P}_{\mathrm{obs}}^{\mathcal{Z}})$, the observational likelihoods coincide for all $n$, so Bayes factors within $\mathcal{E}(\mathbb{P}_{\mathrm{obs}}^{\mathcal{Z}})$ are identically $1$. Consequently,
$\Pi\!\bigl(\mathcal{S}\mid\mathcal{D}_{n}\bigr)\propto\Pi(\mathcal{S})\quad\text{for all }\ \mathcal{S}\in\mathcal{E}(\mathbb{P}_{\mathrm{obs}}^{\mathcal{Z}}),\ \text{all }n.$
We also know that $\Pi(\mathcal{S}|\mathcal{D}_{n})\to 0$ for $S\notin\mathcal{E}(\mathbb{P}_{\mathrm{obs}}^{\mathcal{Z}})$ and $n\to\infty$.
@@
$Q\left(\int\mathbb{P}_{\mathrm{int}}^{\mathcal{S}}\,\Pi(\mathcal{S}\mid\mathcal{D}_{n})\,d\mathcal{S}\right)\;$ $\to\;\int_{\mathcal{E}(\mathbb{P}_{\mathrm{obs}}^{\mathcal{S}_{0}})}Q(\mathbb{P}_{\mathrm{int}}^{\mathcal{S}})\,\Pi(\mathcal{S})\,d\mathcal{S}$ (13)
$=w_{0}\,\int_{\mathcal{Z}}Q(\mathbb{P}_{\mathrm{int}}^{\mathcal{S}})\,\Pi(\mathcal{S})\,d\mathcal{S}\;+\;(1-w_{0})\,Q\!\bigl(\mathbb{P}_{\mathrm{int}}^{\mathcal{W}}\bigr),$ (14)
$\neq Q\!\bigl(\mathbb{P}_{\mathrm{int}}^{\mathcal{W}}\bigr),$ (15)
which shows that $\Pi$ is not well-specified for $\mathcal{C}$. ∎
@@ D.3 Implementation details of baselines
For the standard CATE setting baselines, we follow the implementation from https://github.com/AliciaCurth/CATENets/tree/main for most of the CATE estimators, including S-learner (Künzel et al., 2019), T-learner (Künzel et al., 2019), TARNet (Shalit et al., 2017b), X-leaner (Künzel et al., 2019), DR-learner (Kennedy, 2023b), RA-learner (Curth & van der Schaar, 2021). For the foundation model baselines, we follow the author implementation from https://github.com/vdblm/CausalPFN/tree/main for CausalPFN (Balazadeh et al., 2025); we follow the author implementation from https://github.com/jr2021/Do-PFN for DoPFN (Robertson et al., 2025).
@@ Appendix H Additional experiments
H.1 Additional results on the standard CATE estimation
@@
RA-learner (Curth & van der Schaar, 2021) 0.620 0.421 0.644 0.706 0.523 0.808 0.646 0.353 0.613 0.759
@@
RA-learner (Curth & van der Schaar, 2021) 0.762 $\pm$ 0.14
@@
H.2 Results on other datasets
@@
H.3 Additional results for the IV setting
@@
MRIV (Frauen & Feuerriegel, 2022) 0.759 0.632 0.698 1.011 0.348 0.860 0.929 0.707 0.562 0.380
@@
MRIV (Frauen & Feuerriegel, 2022) 0.579 0.631 0.760 0.189 0.586 1.091 0.471 0.880 0.669 0.556

### v3

@@ Appendix A Extended related work
Foundation models are an emerging paradigm that has revolutionized machine learning for various data modalities, particularly language and vision tasks (Devlin, 2018; Lahat et al., 2024; Touvron et al., 2023b; a). The same paradigm is now being explored for tabular data – a modality that underpins the large majority of analyses in science and business (van Breugel and van der Schaar, 2024). Prior-data-fitted networks (PFNs) (Müller et al., 2022) constitute a powerful approach to training tabular foundation models. PFNs are large transformers trained on synthetic data to perform Bayesian inference through in-context learning. TabPFN (Hollmann et al., 2023; 2025) scaled this idea by pairing the transformer with a Bayesian neural network prior over structural causal models (SCMs) and demonstrating state-of-the-art performance on various tabular benchmarks. Subsequent work extended PFNs to time–series forecasting (Hoo et al., 2025) and analyzed their in-context learning abilities theoretically (Nagler, 2023). Critically, all existing PFNs are trained only for predictive tasks and do not target causal estimands; they therefore are not designed for causal inference of treatment effects, which is the goal of our paper.
@@
Causal inference, such as the estimation of average treatment effects, originates from fields like econometrics (Imbens and Angrist, 1994; Angrist, 1990), statistics (van der Laan and Rubin, 2006), and epidemiology (Robins, 1986; 1994). Machine learning methods have been proposed to estimate heterogeneous effects to support personalized decision-making. One line of work are frequentist methods, which often build on semiparametric theory (Robins et al., 1994; Robins, 1999), yielding model-agnostic estimators that are doubly robust and Neyman-orthogonal (van der Laan, 2006; Chernozhukov et al., 2018; Nie and Wager, 2021; Foster and Syrgkanis, 2023; Kennedy, 2023a). Another line of work builds upon specific machine learning methods/ architectures such as regression trees (Wager and Athey, 2018) or neural networks (Johansson et al., 2016; Shalit et al., 2017a; Shi et al., 2019) and adopts them to causal inference. Bayesian alternatives include Bayesian additive regression trees (Hahn et al., 2020) or Gaussian-process counterfactual regression (Alaa and van der Schaar, 2017). However, all of the existing estimators above must be retrained for every new dataset. In contrast, our CausalFM allows for pre-trained models to approximate Bayesian causal inference.
A.1 Differences between CausalFM, CausalPFN, and Do-PFN
Identifiability. CausalFM separates identifiability from estimation. The central motivation of CausalFM is that causal identification (choosing an identifiable setting such as back-door, IV, or front-door) must be handled before estimation. This mirrors classical causal inference practice and ensures that the PFN only learns within an identifiable causal setting. Our reasoning for identifiability is as follows:
(1) Asymptotically unbiased causal inference: As we show in Theorem 4.3, incorporating identifiability assumptions into the prior is necessary for asymptotically unbiased causal inference. As a consequence, methods that ignore identifiability assumptions yield biased causal effect estimates, even if we collect large amounts of data. This is highly undesirable in practice.
(2) Informative predictive-posterior distributions: If we do not impose any assumption on the DGP, it is well known that causal effect estimation is not just fundamentally biased, but also that this bias can be of arbitrary size. For example, the backdoor-adjustment bias due to omitted unobserved confounding can be written in closed form depending on confounding strength. Thus, if the PFN-prior assigns positive probability mass for DGPs with arbitrary confounding strength, the predictive-posterior must respect the possibility of arbitrarily biased treatment effects, thus rendering PFN-based inference completely noninformative.
(3) Clear separation between domain knowledge and statistical inference: One might argue that a possible remedy would be to restrict the PFN prior only to DGPs with somewhat “weak” identifiability violations (e.g., weak unobserved confounding). However, we argue that this would correspond to assumptions/ domain knowledge on the DGP, similar to those in our paper, that must be made transparent for practitioners and could also possibly be violated.
In short, our paper follows established causal inference philosophy and separates identifiability from estimation: the identifiability step (choosing the causal setting) requires careful modeling and usage of domain knowledge, while the estimation step can be handed over to our CausalFM. If practitioners suspect identifiability assumptions may be violated, we recommend performing causal sensitivity analysis to assess the extent of potential violations.
In contrast, CausalPFN implicitly assumes only back-door adjustment and therefore cannot handle IV or front-door, resulting in bias under unobserved confounding. Do-PFN mixes many causal graphs in a single prior without conditioning on which setting is identifiable, which, as our Theorem 4.3 shows, can lead to asymptotically biased estimates and non-informative posteriors.
Prior construction. This philosophy requires a fundamentally different prior construction. Because identifiability is encoded at the level of causal structure, CausalFM introduces C-SCM priors and C-DAGs that enforce the assumptions required by each identifiability strategy. Do-PFN does not encode identifiability constraints into the prior family for their prior construction. CausalPFN is restricted solely to back-door adjustment.
Theoretical guarantees. Beyond this framework, we contribute new theoretical results showing that identifiability must be incorporated into PFN priors. Theorem 4.3 proves that if a PFN prior places nonzero mass on SCMs that violate the identifiability conditions of the chosen setting, then the resulting posterior predictive interventional distribution is necessarily misspecified and cannot yield consistent causal effect estimates, even with infinite data. This explains the empirical behavior of Do-PFN, which may return non-informative posteriors when its prior includes SCMs with strong unobserved confounding and no valid instruments. Our theoretical results show that this issue is structural, not merely an implementation detail.
Empirical performance. Empirically, our model outperforms others in different settings. Besides, we also have experiments showing the necessity to have the correct identifiability assumption in the prior specification.
@@ Appendix C Proofs
Assume, for contradiction, that $\Pi(\mathcal{Z})>0$. Then, there exists an observational distribution $P_{\mathrm{obs}}\in\mathcal{P}_{\mathrm{obs}}$ such that
$w:=\Pi\!\bigl(\mathcal{Z}\cap\mathcal{E}(P_{\mathrm{obs}})\bigr)>0,\qquad\mathcal{E}(P_{\mathrm{obs}}):=\{\mathcal{S}:\mathbb{P}_{\mathrm{obs}}^{\mathcal{S}}=P_{\mathrm{obs}}\}.$ (12)
Let $\mathcal{W}:=\mathcal{E}(P_{\mathrm{obs}})\setminus\mathcal{Z}$ denote the non-violating subset of the same observational equivalence class. Pick any $\mathcal{S}^{w}\in\mathcal{W}$ with $\Pi(\mathcal{S}^{w})>0$ (such an SCM exists if $\Pi$ is well-specified for $\mathcal{C}$ on $P_{\mathrm{obs}}$), and draw data $\mathcal{D}_{n}\sim\mathbb{P}_{\mathrm{obs}}^{\mathcal{S}^{w}}=P_{\mathrm{obs}}$.
Consider the posterior-predictive interventional functional
$T_{n}\;:=\;Q\!\Bigl(\int\mathbb{P}_{\mathrm{int}}^{\mathcal{S}}\,\Pi(\mathcal{S}\mid\mathcal{D}_{n})\,\mathop{}\!\mathrm{d}\mathcal{S}\Bigr).$ (13)
Because $Q$ is linear, we have
$T_{n}=\int Q\!\bigl(\mathbb{P}_{\mathrm{int}}^{\mathcal{S}}\bigr)\,\Pi(\mathcal{S}\mid\mathcal{D}_{n})\,\mathop{}\!\mathrm{d}\mathcal{S}.$ (14)
Step 1: Split the integral.
Write $\mathcal{R}:=\mathcal{E}(P_{\mathrm{obs}})^{c}$. Then
$T_{n}=\int_{\mathcal{E}(P_{\mathrm{obs}})}Q\!\bigl(\mathbb{P}_{\mathrm{int}}^{\mathcal{S}}\bigr)\,\Pi(\mathcal{S}\mid\mathcal{D}_{n})\,\mathop{}\!\mathrm{d}\mathcal{S}+\int_{\mathcal{R}}Q\!\bigl(\mathbb{P}_{\mathrm{int}}^{\mathcal{S}}\bigr)\,\Pi(\mathcal{S}\mid\mathcal{D}_{n})\,\mathop{}\!\mathrm{d}\mathcal{S}.$ (15)
Step 2: Posterior concentrates on the observational equivalence class.
Under standard Bayesian consistency conditions for the observational model (e.g., KL support of $\Pi$ at $P_{\mathrm{obs}}$), we have
$\Pi\!\bigl(\mathcal{R}\mid\mathcal{D}_{n}\bigr)\to 0,\qquad n\to\infty,$ (16)
and, hence, the last integral over $\mathcal{R}$ vanishes asymptotically (assuming the usual integrability for $Q$).
Step 3: Inside $\mathcal{E}(P_{\mathrm{obs}})$, the posterior is proportional to the prior.
For every $\mathcal{S}\in\mathcal{E}(P_{\mathrm{obs}})$, the observational likelihood is identical because $\mathbb{P}_{\mathrm{obs}}^{\mathcal{S}}=P_{\mathrm{obs}}$. Therefore, for any measurable $A\subseteq\mathcal{E}(P_{\mathrm{obs}})$ and all $n$, we yield
$\Pi(A\mid\mathcal{D}_{n})=\frac{\Pi(A)}{\Pi(\mathcal{E}(P_{\mathrm{obs}}))}\qquad\text{and}\qquad\Pi(\mathop{}\!\mathrm{d}\mathcal{S}\mid\mathcal{D}_{n},\ \mathcal{S}\in\mathcal{E}(P_{\mathrm{obs}}))=\frac{\Pi(\mathop{}\!\mathrm{d}\mathcal{S})}{\Pi(\mathcal{E}(P_{\mathrm{obs}}))}.$ (17)
Combining with Step 2 yields
$T_{n}\;\longrightarrow\;\frac{1}{\Pi(\mathcal{E}(P_{\mathrm{obs}}))}\int_{\mathcal{E}(P_{\mathrm{obs}})}Q\!\bigl(\mathbb{P}_{\mathrm{int}}^{\mathcal{S}}\bigr)\,\Pi(\mathop{}\!\mathrm{d}\mathcal{S}),\qquad n\to\infty.$ (18)
Step 4: The limit cannot equal the identifiable target.
Split the integral over $\mathcal{E}(P_{\mathrm{obs}})$ into $\mathcal{W}$ and $\mathcal{Z}\cap\mathcal{E}(P_{\mathrm{obs}})$:
$\int_{\mathcal{E}(P_{\mathrm{obs}})}Q\!\bigl(\mathbb{P}_{\mathrm{int}}^{\mathcal{S}}\bigr)\,\Pi(\mathop{}\!\mathrm{d}\mathcal{S})=\int_{\mathcal{W}}Q\!\bigl(\mathbb{P}_{\mathrm{int}}^{\mathcal{S}}\bigr)\,\Pi(\mathop{}\!\mathrm{d}\mathcal{S})+\int_{\mathcal{Z}\cap\mathcal{E}(P_{\mathrm{obs}})}Q\!\bigl(\mathbb{P}_{\mathrm{int}}^{\mathcal{S}}\bigr)\,\Pi(\mathop{}\!\mathrm{d}\mathcal{S}).$ (19)
By definition of $\mathcal{W}$ and identifiability, for every $\mathcal{S}\in\mathcal{W}$ we have $Q(\mathbb{P}_{\mathrm{int}}^{\mathcal{S}})=\bar{Q}(P_{\mathrm{obs}})$, and thus
$\int_{\mathcal{W}}Q\!\bigl(\mathbb{P}_{\mathrm{int}}^{\mathcal{S}}\bigr)\,\Pi(\mathop{}\!\mathrm{d}\mathcal{S})=\bar{Q}(P_{\mathrm{obs}})\,\Pi(\mathcal{W}).$ (20)
@@
$\lim_{n\to\infty}T_{n}=\frac{\bar{Q}(P_{\mathrm{obs}})\,\Pi(\mathcal{W})+\int_{\mathcal{Z}\cap\mathcal{E}(P_{\mathrm{obs}})}Q\!\bigl(\mathbb{P}_{\mathrm{int}}^{\mathcal{S}}\bigr)\,\Pi(\mathop{}\!\mathrm{d}\mathcal{S})}{\Pi(\mathcal{W})+\Pi\!\bigl(\mathcal{Z}\cap\mathcal{E}(P_{\mathrm{obs}})\bigr)}.$ (21)
By the non-cancellation assumption (applied to this $P_{\mathrm{obs}}$) and $w>0$, this limit is not equal to $\bar{Q}(P_{\mathrm{obs}})$.
Finally, since $\mathcal{D}_{n}\sim\mathbb{P}_{\mathrm{obs}}^{\mathcal{S}^{w}}$ and $\mathcal{C}$ is identifiable,
$Q\!\bigl(\mathbb{P}_{\mathrm{int}}^{\mathcal{S}^{w}}\bigr)=\bar{Q}(P_{\mathrm{obs}}).$ (22)
Thus $\lim_{n\to\infty}T_{n}\neq Q(\mathbb{P}_{\mathrm{int}}^{\mathcal{S}^{w}})$, contradicting that $\Pi$ is well-specified for $\mathcal{C}$. Therefore, it must be that $\Pi(\mathcal{Z})=0$. ∎
@@ D.1 Implementation Details of data prior
Input format. CausalFM operates as an in-context learner like other foundation models, meaning it approximates Bayesian inference by conditioning on a dataset provided in its context window. Therefore, it requires an entire dataset to make predictions for specific query samples. (1) Input structure: The model accepts a dataset $\mathcal{D}_{n}=\left\{\left(x_{i},a_{i},y_{i}\right)\right\}_{i=1}^{n}$ acting as the context (or support set) and a query point $x_{\text{query }}$ (or a batch of query points). (2) Mechanism: The transformer processes the entire sequence of observed data $\mathcal{D}_{n}$ using self-attention to extract context-dependent representations. It then outputs the posterior predictive distribution for the causal quantity (e.g., CATE given the context $\mathcal{D}_{n}$ and the specific query $x_{\text{query }}$. (3) Comparison to fine-tuning baselines: In practice, standard baselines require an explicit training phase on a training set before evaluation. In contrast, CausalFM takes the “training” data as the input context (support set) and directly generates predictions for the test data (query set) in a single forward pass.
@@ D.3 Implementation details of baselines
For the standard CATE setting baselines, we follow the implementation from https://github.com/AliciaCurth/CATENets/tree/main for most of the CATE estimators, including S-learner (Künzel et al., 2019), T-learner (Künzel et al., 2019), TARNet (Shalit et al., 2017b), X-leaner (Künzel et al., 2019), DR-learner (Kennedy, 2023b), RA-learner (Curth and van der Schaar, 2021b). For the foundation model baselines, we follow the author implementation from https://github.com/vdblm/CausalPFN/tree/main for CausalPFN (Balazadeh et al., 2025); we follow the author implementation from https://github.com/jr2021/Do-PFN for DoPFN (Robertson et al., 2025).
@@ Appendix H Additional experiments
H.1 Evaluation in the front-door adjustment setting
H.1.1 Baselines for front-door adjustment setting
In contrast to the standard CATE or IV settings, there are few established baselines for the front-door case. Identification in this setting is enabled through Pearl’s front-door formula (Pearl, 2009). The natural baseline is therefore the plug-in front-door learner, which estimates the necessary nuisance components, i.e., $P(M\mid A,X)$, $P(A\mid X)$, and $\mathbb{E}[Y\mid M,X]$ and substitutes them into the identification formula to recover causal quantities. To assess the role of model flexibility in estimating these nuisance functions, we implement the plug-in learner with different regression methods, including linear regression, Random Forests, and neural networks.
H.1.2 Results for front-door adjustment setting
Table 6.2 reports the averaged PEHE across datasets. We observe that CausalFM achieves competitive CATE estimation. Importantly, these results hold without requiring model retraining for our model, demonstrating the adaptability of our approach to the front-door setting.
H.2 Additional results on the standard CATE estimation
@@
RA-learner (Curth and van der Schaar, 2021b) 0.620 0.421 0.644 0.706 0.523 0.808 0.646 0.353 0.613 0.759
@@
RA-learner (Curth and van der Schaar, 2021b) 0.762 $\pm$ 0.14
@@
H.3 Results on other datasets
@@
H.4 Additional results for the IV setting
@@
MRIV (Frauen and Feuerriegel, 2022) 0.759 0.632 0.698 1.011 0.348 0.860 0.929 0.707 0.562 0.380
@@
MRIV (Frauen and Feuerriegel, 2022) 0.579 0.631 0.760 0.189 0.586 1.091 0.471 0.880 0.669 0.556
@@
H.5 Additional results for misspecification and prior design choices
Figure 2: Robustness of our model to difference choices of the prior.
Misspecification of causal settings. We conduct experiments to study the sensitivity to using an incorrect identifiability strategy. Specifically, we generate data from (i) an IV SCM and (ii) a front-door SCM (as in our main experiments), and compare a model trained under the correct identifiability design (IV or front-door, respectively) with a model trained under an incorrect back-door design. The results are in Table 9. As expected, using a misspecified identifiability strategy consistently worsens the PEHE. This highlights the importance of including the correct identifiability assumption into the prior specification, which we propose for CausalFM.
Table 9: Analysis of the effect of misspecified identifiability strategies.
Data Generating SCM Strategy Identifiability Used PEHE
IV SCM Correct IV 0.422
Incorrect Back-door 0.489
Front-door SCM Correct Front-door 0.847
Incorrect Back-door 0.876
Prior design choices. We also analyze the robustness of our model to different choices of the prior. In particular, we vary the strength of unobserved confounding in the data-generating process, controlled by a parameter $\alpha\in[0,1]$. The results in Fig. 2 show that our model remains robust as $\alpha$ increases. This again confirms the strong performance of CausalFM.
H.6 Computational time
We report computation time in this section. For our method and other foundation-model-based approaches, we show inference time since these models do not need fine-tuning after pretraining. For standard baselines, which must be trained for each dataset separately, we report the average total time per dataset, including both training and inference. As shown in Tables H.6 and H.6, our model is highly efficient.
Table 10: Overall time comparison for standard CATE setting.
Method Time (s)
[…]
S-learner (Künzel et al., 2019) $2.76\times 10^{0}$
T-learner (Künzel et al., 2019) $3.21\times 10^{0}$
TARNet (Shalit et al., 2017b) $3.98\times 10^{0}$
DR-learner (Kennedy, 2023b) $1.78\times 10^{1}$
RA-learner (Curth and van der Schaar, 2021b) $1.24\times 10^{1}$
X-learner (Künzel et al., 2019) $1.93\times 10^{1}$
[…]
CausalPFN (Balazadeh et al., 2025) $1.27\times 10^{0}$
DoPFN (Robertson et al., 2025) $2.31\times 10^{0}$
CausalFM (ours) $4.90\times 10^{-1}$
Lower = better. Reported: Time in seconds.
Table 11: Overall time comparison for IV setting.
Method Time (s)
[…]
KIV (Singh et al., 2019) $3.24\times 10^{-1}$
DRIV (Syrgkanis et al., 2019) $3.87\times 10^{1}$
DeepIV (Hartford et al., 2017) $1.27\times 10^{1}$
DeepGMM (Bennett et al., 2019) $1.85\times 10^{1}$
DMLIV (Syrgkanis et al., 2019) $1.85\times 10^{1}$
DFIV (Xu et al., 2021) $1.74\times 10^{1}$
MRIV (Frauen and Feuerriegel, 2022) $1.56\times 10^{1}$
[…]
DoPFN (Robertson et al., 2025) $6.53\times 10^{0}$
CausalFM (ours) $4.72\times 10^{-1}$
Lower = better. Reported: Time in seconds.